# arXiv:2602.13749v3[cond-mat.str-el] 02 Mar 2026

### v3

@@ VI.1 Non-Hermitian NRG results for the PT symmetric model
Figure 6: (a) The nHNRG results about the thermodynamic quantities for $\mathcal{PT}$-symmetric $n=2$ NHMCK. (a) shows the real part of the energy spectrum, where the dark dot, blue, green, and red curves are the results for $\theta=0,\ 0.1,\ 0.5,\ 1.0$, respectively. On the right side, we show the energy levels and the degeneracies (in the bracket) predicted by BCFT around the low-energy fixed point, which is in exact agreement with the NRG results. (b) shows the partition function as a function of temperature for various $\theta$, which is positive for all energy scales. (c) The calculated impurity entanglement entropy $S_{\rm imp}$ and impurity specific heat $C_{\rm imp}$ as a function of temperature. (d) and (e) are the zoom-in plot of (c) in the intermediate temperature range, which displays non-negligible shifts of thermodynamic quantities due to the non-Hermitian effects.
In Fig.6(b), we further observe that the partition function always maintains positive for the $\mathcal{PT}$-symmetric model, even if the $\mathcal{PT}$ symmetry can be spontaneously broken. In Fig.6(c), we plot the thermodynamic properties associated with the impurity. As shown, both the impurity entropy $S_{\rm imp}$ and the impurity specific heat $C_{\rm imp}$ flow to the strong-coupling overscreening fixed point at low temperatures for various $\theta$. Fig.6(d) and (e) show the zoom-in data of $S_{\rm imp}$ and $C_{\rm imp}$ with taking more different values of $\theta$, respectively. Interestingly, we find that the thermodynamic quantities displays non-negligible shifts as $\theta$ is enlarged. This indicates that although the non-Hermiticity does not alter the fixed point at low-temperatures, they have nontrivial effects and modify thermodynamic properties at finite temperatures. This will be discussed again in our BCFT calculations in Sec. VII.2 and Appendix D.
VI.2 Applicability of BCFT under PT symmetry
The NHNRG method above has shown that the low-energy spectrum of the $\mathcal{PT}$-symmetric model exactly coincides with the BCFT predictions, i.e., in agreement with the spectrum corresponding to the strongly correlated fixed point. In this section, we will give a more rigorous proof that the low-energy theory of $\mathcal{PT}$-symmetric model is related to the WZW CFT via a similarity transformation. Thus, the BCFT method can be readily applied to study the $\mathcal{PT}$-symmetric NHMCK models, which can further yield, in a non-perturbative way, thermodynamic and transport properties associated with the strong-coupling fixed point, as will be discussed in Sec. VII below.
Let us emphasize again here, for $\mathcal{PT}$ symmetric systems involving no exceptional point, the non-Hermitian Hamiltonian $\widetilde{H}$ (we slightly shift our notation here for convenience) can be transformed into a Hermitian Hamiltonian via a similar transformation,
@@ VI.2 Applicability of BCFT under PT symmetry
remains unchanged compared to the Hermition case. Likewise, the Affleck-Ludwig boundary entropy,
@@
also coincides with its Hermitian counterpart. Therefore, we have shown that for $\mathcal{PT}$-symmetric models without exceptional point, the BCFT approach applies directly. In addition, when one calculates thermodynamical quantities in the $\mathcal{PT}$ region, the Boltzmann weight reduces to that of Hermitian case with only a metric inserted in Hilbert space. Therefore, we have shown that if there is no spontaneous $\mathcal{PT}$ symmetry breaking, the BCFT description remains soundly applicable in the low-energy window. Also, this is verified by the consistence with the NHNRG results, where the imaginary energy apparently vanishes in the low-energy spectrum, i.e., Fig.5(a). In addition, the thermodynamic quantities such as specific heat and conductance will be discussed in the following sections and Appendix D.
VII transport properties around fixed points
In this section, we aim to study transport properties associated with the fixed points of the NHMCK model. For the weak-coupling fixed point discussed in Sec. III, since it is well within the description of perturbative theory, we will calculate the corresponding Kondo conductance based on the non-Hermitian Kubo formula. While for the strong-coupling fixed point, which is essentially non-perturbative, we restrict ourselves to the $\mathcal{PT}$-symmetric model and adopt the BCFT approach, whose validity has been justified for the $\mathcal{PT}$ symmetric case in the last section.
VII.1 non-Hermitian Kubo formula approach to the weak-coupling fixed point
@@
which provides a natural generalization of its unitary counterpart. Note that with the denominator, the trace of Eq.(89) is preserved, while the trace of the unnormalized density matrix, $\tilde{\rho}_{t}=e^{-{i}(H_{r}-{i}{H_{i}})t}\rho_{0}e^{{i}(H_{r}+{i}{\mathit{\Gamma}})t}$, which is directly derived from the Lindblad master equation, is generally not 1, unless the Hamiltonian always holds a real energy spectrum. This has been discussed above in Sec. II.
@@
where $H_{K}^{\prime}$ and $H_{K*}$ represent the real part and the imaginary part of the second-order co-tunneling terms, i.e., the Kondo exchange terms. Note that the operator derived from Eq.(90) is Hermitian, and its expectation value is real, owing to Eq.(88) which guarantees the trace condition and thus the particle conservation. This is in consistence with the post-selection on the Lindbalian, i.e., resetting the particle number, which also ensures the trace normalization of the density matrix, resulting the particle number conservation.
After calculating the commutators above, write the current operator is cast into the form,
@@
where $G_{0}$ is the conventional conductance corresponding to the Hermitian case, $G_{*}$ denotes the contribution of imaginary coupling, and $\delta G_{\langle\rangle}$ comes from the norm correction. Further inserting the specific form of $\rho_{r}$, $\rho_{i}$, and $H_{K,*}$, one can readily find that all these three terms are proportional to the square of the exchange coupling around the weak-coupling FP. Specifically, $G_{0}\sim\lambda^{2}_{r}\sim g^{2}_{r}$, and $G_{*}$, $\delta G_{\langle\rangle}\sim\lambda^{2}_{i}\sim g^{2}_{i}$. Further utilizing the RG scaling behavior of $g_{r}$ and $g_{i}$ around the weak-coupling fixed pont, i.e., $g_{r},g_{i}\sim\frac{1}{\ln(T/T_{\mathrm{Kwc}})}$, it is derived that,
@@
Interestingly, this Kondo conductance behavior is distinct from previously reported Kondo systems. Although a similar temperature dependence was found for underscreened Kondo models [84, 85, 86], the conductance therein saturates to finite values for low temperatures. In addition, although zero conductance occurs at zero temperature in the unscreened channel of anisotropic multichannel Kondo models, Fermi-liquid-type behavior, i.e., $(T/T_{K})^{2}$, emerges at low temperatures in that case. Thus, the conductance here is quite unique, and it is attributed to a completely different mechanism–the non-Hermticity induced decoupling of the local moment. The detailed results and comparison of the conductance are shown in Fig.4 of our joint paper [87].
VII.2 BCFT approach to the strong-coupling fixed point
We now study the Kondo conductance for the strong-coupling fixed point using the BCFT approach. Although the perturbative methods fail for this fixed point, we have shown that, when imposing the $\mathcal{PT}$ symmetry, the pseudo-Hermiticity allows us to solve it similarly as the Hermitian stereotype. This leads to the fact that the BCFT results are applicable in this regime, as a deformation of the Hermitian approach.
@@ VII.2 BCFT approach to the strong-coupling fixed point
For the $\mathcal{PT}$ symmetric non-Hermitian Kondo model, however, in the anti-Hermitian sector, a new irrelevant operator that satisfies $\mathcal{PT}$ symmetry can be found, i.e., $i\mathbf{J}_{-1}\phi_{j=2}$. Although the Kac-Moody descendant $\mathbf{J}_{-1}\phi_{j=2}$ itself explicitly breaks $\mathcal{PT}$ symmetry, the imaginary coupling parameter restores the $\mathcal{PT}$-invariance of the whole perturbed Hamiltonian. The non-Hermitian boundary interaction legitimates such anti-Hermitian perturbation terms. This operator has total spin 1 and the scaling dimension $\Delta=2$, and it is the leading order in the anti-Hermitian sector.
@@
However, we mention briefly that, for the $n=2$ NHMCK, the non-Hermitian effect in IR limit is much weaker than the $n=4$ case. This is because the possible leading operators in anti-Hermitian sector therein turns out to be one-level higher descendants (sub-sub-leading terms), e.g., $\sim i\mathbf{J}_{-2}\cdot\phi$, with the scaling dimension $\Delta=5/2$. Nevertheless, this would still lead to non-negligible modifications to the conductance and thermodynamic quantities at finite temperatures. Interestingly, the latter is indeed numerically observed by our non-Hermitian NRG calculations, as shown by Fig.6. Thus, in this aspect, the BCFT and the nHNRG results are again in consistence with each other.
VIII Conclusion and Discussion
The multimethod study presented in this work report the following key findings. First, our developed NHNRG method uncovers numerical signatures of the YSR state beyond the perturbative theory, as a characteristic feature in non-Hermitian Kondo systems. The NRG results presented here are clearly beyond those in Ref. [132], as we further calculate the thermodynamic quantities as a function temperatures. Second, we have shown that non-Hermiticity can generate unconventional Kondo conductance behaviors when the multichannel feature leading to NFL is further considered. Particularly, When NFL is combined with non-Hermiticity, an anomalous upturn would occur in the Kondo conductance around the strong-coupling fixed point, originating from an emergent BCFT with non-Hermitian boundary operators. This is unique for Kondo systems that simultaneously display non-Hermitian and nFL behaviors. For example, for the non-Hermitian Fermi liquid fixed point that occurs for the single-channel case, the impurity is exactly screened, and the LIOs around the strong coupling FP are simply four-fermion operators without imaginary couplings.
We remark that although the Kondo systems are studied in this work, the results here imply that exotic physics could also take place for other non-Hermitian correlated systems without quasi-particles descriptions, for example, the non-Hermitian Luttinger liquid  [145, 146, 147, 148, 149]. Furthermore, regarding the experimental realization, we would to highlight again that cold-atom systems provide an alternative and maybe more feasible platform to realize our predicted Kondo phenomena. Recently, it is reported that the multi-orbital cold atoms can be engineered to host both localized and itinerant degrees of freedom [101, 102], where the dissipation can also be realdiy introduced [103, 37]. Designable dissipation and system-integratable monitoring equipments are also achievable in this platform  [104, 105, 106, 107, 108, 109]. Last, since we restrict ourselves to $\mathcal{PT}$-symmetric case when calculating the Kondo conductance around the strong-coupling fixed point, it is important to loosen the $\mathcal{PT}$ restriction and investigate how the conductance (and also the other thermodynamic quantities) would be modified by advanced methods, such as Ref. [112]. This points to an interesting direction, i.e., the symmetry-resolved non-Hermitian effects in correlated systems.
@@ Acknowledgments
From renormalized QD Green’s function, it is known that a imaginary chemical potential term is generated.
@@
Thus, the Hamiltonian can be written in matrix form, and the entries $\mathcal{H}_{ss^{\prime}}$, $\mathcal{H}_{dd^{\prime}}$ and $H_{sd}$ describe the system, quantum dot and cross term respectively.
@@
In this Appendix, we present more details in determining the critical line separating the strong- and weak-coupling fixed point in the perturbative RG theory. Starting from the RG flow in Eq.(22) and integrating the energy cutoff from $\Lambda_{0}$ to $\Lambda$, we obtain
@@
Appendix D Thermodynamics analyzed by BCFT
In this appendix section, we will give more BCFT details on the thermodynamics of $\mathcal{PT}$-symmetric model.

### v4

@@ VII.1 Non-Hermitian NRG results for the PT-symmetric model
Figure 6: (a) The NHNRG results of the thermodynamic quantities for $\mathcal{PT}$-symmetric $n=2$ NHMCK. (a) shows the real part of the energy spectrum, where the dark dot, blue, green, and red curves are the results for $\theta=0,\ 0.1,\ 0.5,\ 1.0$, respectively. On the right side, we show the energy levels and the degeneracies (in the bracket) predicted by BCFT around the low-energy fixed point, which is in exact agreement with the NRG results. (b) shows the partition function as a function of temperature for various $\theta$, which is positive for all energy scales. (c) The calculated impurity entanglement entropy $S_{\rm imp}$ and impurity specific heat $C_{\rm imp}$ as a function of temperature. (d) and (e) are the zoom-in plot of (c) in the intermediate temperature range, which displays non-negligible shifts of thermodynamic quantities due to the non-Hermitian effects.
In Fig.6(b), we further observe that the partition function always maintains positive for the $\mathcal{PT}$-symmetric model, even if the $\mathcal{PT}$ symmetry can be spontaneously broken. In Fig.6(c), we plot the thermodynamic properties associated with the impurity. As shown, both the impurity entropy $S_{\rm imp}$ and the impurity specific heat $C_{\rm imp}$ flow to the strong-coupling overscreening fixed point at low temperatures for various $\theta$. Fig.6(d) and (e) show the zoom-in data of $S_{\rm imp}$ and $C_{\rm imp}$ with taking more different values of $\theta$, respectively. Interestingly, we find that the thermodynamic quantities display non-negligible shifts as $\theta$ is enlarged. This indicates that although the non-Hermiticity does not alter the fixed point at low-temperatures, it has nontrivial effects and modify thermodynamic properties at finite temperatures. This will be discussed again in our BCFT calculations in Sec. VIII.2 and Appendix E.
VII.2 Applicability of BCFT under PT symmetry
The NHNRG method above has shown that the low-energy spectrum of the $\mathcal{PT}$-symmetric model exactly coincides with the BCFT predictions, i.e., in agreement with the spectrum corresponding to the strongly correlated fixed point. In this section, we will give a more rigorous proof that the low-energy theory of $\mathcal{PT}$-symmetric model is related to the Wess-Zumino-Witten (WZW) CFT via a similarity transformation. Thus, the BCFT method can be readily applied to study the $\mathcal{PT}$-symmetric NHMCK models, which can further yield, in a non-perturbative way, thermodynamic and transport properties associated with the strong-coupling fixed point, as will be discussed in Sec. VIII below.
Let us emphasize again, for $\mathcal{PT}$ symmetric systems involving no exceptional point, the non-Hermitian Hamiltonian $\widetilde{H}$ (we slightly shift our notation here for convenience) can be transformed into a Hermitian Hamiltonian via a similarity transformation,
@@ VII.2 Applicability of BCFT under PT symmetry
remains unchanged compared to the Hermitian case. Likewise, the Affleck-Ludwig boundary entropy,
@@
also coincides with its Hermitian counterpart. Therefore, we have shown that for $\mathcal{PT}$-symmetric models without exceptional point, the BCFT approach applies directly. In addition, when one calculates thermodynamical quantities in the $\mathcal{PT}$ region, the Boltzmann weight reduces to that of Hermitian case with only a metric inserted in Hilbert space. Therefore, we have shown that if there is no spontaneous $\mathcal{PT}$ symmetry breaking, the BCFT description remains soundly applicable in the low-energy window. Also, this is verified by the consistency with the NHNRG results, where the imaginary energy apparently vanishes in the low-energy spectrum, i.e., Fig.5(a). In addition, the thermodynamic quantities such as specific heat and conductance will be discussed in the following sections and Appendix E.
VIII transport properties around fixed points
In this section, we aim to study the transport properties associated with the fixed points of the NHMCK model. For the weak-coupling fixed point discussed in Sec. IV, since it is well within the description of perturbation theory, we will calculate the corresponding Kondo conductance based on the non-Hermitian Kubo formula. While for the strong-coupling fixed point, which is essentially non-perturbative, we restrict ourselves to the $\mathcal{PT}$-symmetric model and adopt the BCFT approach, whose validity has been justified for the $\mathcal{PT}$ symmetric case in the previous section.
VIII.1 non-Hermitian Kubo formula approach to the weak-coupling fixed point
@@
which provides a natural generalization of its unitary counterpart. Note that with the denominator, the trace of Eq.(91) is preserved, while the trace of the unnormalized density matrix, $\tilde{\rho}_{t}=e^{-{i}(H_{r}-{i}{H_{i}})t}\rho_{0}e^{{i}(H_{r}+{i}{\mathit{\Gamma}})t}$, which is directly derived from the Lindblad master equation, is generally not 1, unless the Hamiltonian always holds a real energy spectrum. This has been discussed above in Sec. III.
@@
where $H_{K}^{\prime}$ and $H_{K*}$ represent the real part and the imaginary part of the second-order co-tunneling terms, i.e., the Kondo exchange terms. Note that the operator derived from Eq.(92) is Hermitian, and its expectation value is real, owing to Eq.(90) which guarantees the trace condition and thus the particle conservation. This is consistent with the post-selection on the Lindbladian, i.e., resetting the particle number, which also ensures the trace normalization of the density matrix, resulting in the particle number conservation.
After calculating the commutators above, the current operator is cast into the form,
@@
where $G_{0}$ is the conventional conductance corresponding to the Hermitian case, $G_{*}$ denotes the contribution of imaginary coupling, and $\delta G_{\langle\rangle}$ comes from the norm correction. Further inserting the specific form of $\rho_{r}$, $\rho_{i}$, and $H_{K,*}$, one can readily find that all these three terms are proportional to the square of the exchange coupling around the weak-coupling FP. Specifically, $G_{0}\sim\lambda^{2}_{r}\sim g^{2}_{r}$, and $G_{*}$, $\delta G_{\langle\rangle}\sim\lambda^{2}_{i}\sim g^{2}_{i}$. Further utilizing the RG scaling behavior of $g_{r}$ and $g_{i}$ around the weak-coupling fixed point, i.e., $g_{r},g_{i}\sim\frac{1}{\ln(T/T_{\mathrm{Kwc}})}$, it is derived that,
@@
Interestingly, this Kondo conductance behavior is distinct from previously reported Kondo systems. Although a similar temperature dependence was found for underscreened Kondo models [85, 86, 87], the conductance therein saturates to finite values for low temperatures. In addition, although zero conductance occurs at zero temperature in the unscreened channel of anisotropic multichannel Kondo models, Fermi-liquid-type behavior, i.e., $(T/T_{K})^{2}$, emerges at low temperatures in that case. Thus, the conductance here is quite unique, and it is attributed to a completely different mechanism–the non-Hermiticity induced decoupling of the local moment. The detailed results and comparison of the conductance are shown in Fig.4 of our joint paper [88].
VIII.2 BCFT approach to the strong-coupling fixed point
We now study the Kondo conductance for the strong-coupling fixed point using the BCFT approach. Although the perturbative methods fail for this fixed point, we have shown that, when imposing the $\mathcal{PT}$ symmetry, the pseudo-Hermiticity allows us to solve it in a similar manner as the Hermitian stereotype. This leads to the fact that the BCFT results are applicable in this regime, as a deformation of the Hermitian approach.
@@ VIII.2 BCFT approach to the strong-coupling fixed point
For the $\mathcal{PT}$ symmetric non-Hermitian Kondo model, however, in the anti-Hermitian sector, a new irrelevant operator that satisfies $\mathcal{PT}$ symmetry can be found, i.e., $i\mathbf{J}_{-1}\phi_{j=2}$. Although the Kac-Moody descendant $\mathbf{J}_{-1}\phi_{j=2}$ itself explicitly breaks $\mathcal{PT}$ symmetry, the imaginary coupling parameter restores the $\mathcal{PT}$-invariance of the whole perturbed Hamiltonian. The non-Hermitian boundary interaction legitimizes such anti-Hermitian perturbation terms. This operator has total spin 1 and the scaling dimension $\Delta=2$, and it is the leading order in the anti-Hermitian sector.
@@
However, we mention briefly that, for the $n=2$ NHMCK, the non-Hermitian effect in IR limit is much weaker than the $n=4$ case. This is because the possible leading operators in anti-Hermitian sector therein turn out to be one-level higher descendants (sub-sub-leading terms), e.g., $\sim i\mathbf{J}_{-2}\cdot\phi$, with the scaling dimension $\Delta=5/2$. Nevertheless, this would still lead to non-negligible modifications to the conductance and thermodynamic quantities at finite temperatures. Interestingly, the latter is indeed numerically observed by our non-Hermitian NRG calculations, as shown by Fig.6. Thus, in this aspect, the BCFT and the NHNRG results are again in consistency with each other.
IX Conclusion and Discussion
We remark that although the Kondo systems are studied in this work, the results here imply that exotic physics could also take place for other non-Hermitian correlated systems without quasi-particle descriptions, for example, the non-Hermitian Luttinger liquid  [146, 147, 148, 149, 150]. Furthermore, regarding the experimental realization, we would like to highlight again that cold-atom systems provide an alternative and maybe more feasible platform to realize our predicted Kondo phenomena. Recently, it is reported that the multi-orbital cold atoms can be engineered to host both localized and itinerant degrees of freedom [103, 104], where the dissipation can also be readily introduced [105, 37]. Designable dissipation and system-integrable monitoring equipment are also achievable in this platform  [106, 107, 108, 109, 110, 111]. Last, since we restrict ourselves to $\mathcal{PT}$-symmetric case when calculating the Kondo conductance around the strong-coupling fixed point, it is important to loosen the $\mathcal{PT}$ restriction and investigate how the conductance (and also the other thermodynamic quantities) would be modified by advanced methods, such as Ref. [114]. This points to an interesting direction, i.e., the symmetry-resolved non-Hermitian effects in correlated systems.
@@ Acknowledgments
From renormalized QD Green’s function, it is known that an imaginary chemical potential term is generated.
@@
Thus, the Hamiltonian can be written in matrix form, and the entries $\mathcal{H}_{ss^{\prime}}$, $\mathcal{H}_{dd^{\prime}}$ and $H_{sd}$ describes the system, quantum dot and cross term respectively.
@@
In this Appendix, we present more details of determining the critical line separating the strong- and weak-coupling fixed point in the perturbative RG theory. Starting from the RG flow in Eq.(22) and integrating the energy cutoff from $\Lambda_{0}$ to $\Lambda$, we obtain
@@
Appendix D Convergence analysis of NHNRG method
Figure 7: Convergence analysis of the NHNRG method. Impurity entropy $S_{\rm imp}$ is calculated with different values of scaling parameter $\Lambda$ and different truncation schemes. (a-c) are results obtained using the “LowReMag”, “LowIm”, and “LowMag” schemes respectively, and $\Lambda=6$. (d-f) show the ‘LowRe” results for $\Lambda=4$, 6, 8, respectively. The data in panel (e) are extracted from Fig. 6(c) in the main text, and serves as a reference here. The number of kept states is $N_{\rm k}=2048$ for all panels.
In this appendix, we perform a convergence analysis of our numerical method. Taking the $\mathcal{PT}$-symmetric two-channel Kondo model as an example, we calculate the impurity entropy $S_{\rm imp}$ using different truncation schemes (Fig.7(a)-(c)) and different values of the scaling parameter $\Lambda$ (Fig.7(d)-(f)). Results in Fig. 7(e) are extracted from Fig. 6(c), which are obtained with the “LowRe” truncation scheme and $\Lambda=6$ , and can be used as a reference to compare with other numerical settings.
In addition to the “LowRe” scheme mentioned in the main text, we have testified three other trunction schemes, as were originally introduced in Ref. [84]. In the “LowIm” scheme, eigenenergies are sorted primarily by their imaginary parts and secondarily by their real parts. In the “LowMag” scheme, eigenenergies are sorted simply by their magnitudes. In the hybrid “LowReMag” scheme, one firstly searches for the ground state energy within the “LowRe” scheme, and then sorts the excited eigenenergies using the “LowMag” scheme.
We compare the results obtained from different truncation schemes for a fixed scaling parameter with $\Lambda=6$. As shown, the data obtained with the “LowReMag” scheme (Fig. 7(a)) are well converged and agree with those from the “LowRe” scheme employed in the main text (Fig. 6(c) and Fig. 7(e)). However, under the “LowIm” scheme (Fig. 7(b)), the impurity entropy fails to approach towards the common fixed points for all values of $\theta$. The values of $S_{\rm imp}$ at low temperatures exhibit randomness and cannot even be reduced to the Hermitian result for $\theta=0.00$. Similarly, under the “LowMag” scheme (Fig. 7(c)), the system flows to a new fixed point, where the infrared impurity entropy saturates at $S_{\rm imp}(T\to 0)\approx 5.5{\rm ln}2$, which is much greater than the local moment value. Thus, the “LowIm” and “LowMag” scheme either lack the numerical convergence or fail to produce reliable physical results.
The above results are in consistence with Ref. [84], which remarks that both the “LowRe” and “LowReMag” schemes perform well and are numerically stable, whereas the “LowIm” and “LowMag” schemes fail to produce reliable outcomes.
In Fig. 7(d)-(f), we show the impurity entropy obtained using the “LowRe” scheme for different scaling parameters, $\Lambda=4,\ 6,\ 8$. It can be found the results are not affected by the choice of $\Lambda$, which only displays minor quantitative difference at high temperatures. This aligns with the conventional Hermitian NRG method.
Appendix E Thermodynamics analyzed by BCFT
In this appendix section, we will give more BCFT details of the thermodynamics of $\mathcal{PT}$-symmetric model.